Opened 10 years ago

Closed 9 years ago

Last modified 2 years ago

#229 closed flight processing (fixed)

GB08/08, flight day 075b/2009, Wytham Woods

Reported by: benj Owned by: knpa
Priority: alpha 4 high Milestone: 2009 Data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by benj)

Data location: ~arsf/arsf_data/2009/flight_data/uk/GB08_08-2009_075b_Wytham_Woods

Data arrived from ARSF via SATA J on 19th March 2009.

Scientific details: see /users/rsg/arsf/arsf_data/2008/ARSF_Applications/GB_and_NL/GB08-08 Proposal and Science Case.pdf

Priority: a4H

PI: R. Harding

Sensors:

  • Eagle
  • Hawk
  • ALS-II Lidar

Quicklook

TBD.

Change History (27)

comment:1 Changed 10 years ago by benj

  • Description modified (diff)

comment:2 Changed 10 years ago by benj

  • Summary changed from GB08/08, flight day 075b, Wytham Woods to GB08/08, flight day 075b/2009, Wytham Woods

comment:3 Changed 10 years ago by benj

  • Description modified (diff)
  • Priority changed from medium to high

comment:4 Changed 10 years ago by jatt

  • Owner set to jatt

comment:5 Changed 10 years ago by mark1

  • Owner changed from jatt to mark1

comment:6 Changed 10 years ago by mark1

GPT Timing offsets:

Eagle 1: -0.015

Eagle 2: -0.01

Eagle 3: -0.05

Eagle 4: -0.07

Hawk 1: -0.04

Hawk 2: -0.01

Hawk 3: -0.05

Hawk 4: -0.1

comment:7 Changed 10 years ago by mark1

Due to issues with Eagle/Hawk sensor misalignment (see #220), have performed per line boresighting. Boresight values used to process data: (P,R,H)

Eagle line 1: -0.113 -0.7387 0.360

Eagle line 2: -0.256 -0.3751 0.370

Eagle line 3: -0.110 -0.6877 0.537

Eagle line 4: 0.019 -0.4037 0.536

Hawk line 1: -0.06 0.15 0.3

Hawk line 2: -0.263 0.297 0.475

Hawk line 3: -0.103 0.19 0.527

Hawk line 4: 0.108 0.266 0.557

comment:8 Changed 10 years ago by mark1

  • Owner changed from mark1 to jatt

Delivery directory created at

~mark1/scratch_space/2009_data/GB08_08-2009_075b_Wytham_Woods/delivery/20090423/GB08_08/

still awaiting final quality report

comment:9 Changed 10 years ago by mark1

up to date data quality report added

comment:10 Changed 10 years ago by jatt

all looks good, rsynced to: /users/rsg/arsf/arsf_data/in_progress/2009/flight_data/uk/GB08_08-2009_075b_Wytham_Woods

ready for delivery

comment:11 Changed 10 years ago by jatt

  • Owner changed from jatt to mark1

Eagle and hawk delivered 27/04/2009

comment:12 Changed 10 years ago by mggr

  • Milestone set to 2009 Data processing completion
  • Priority changed from alpha 5 to alpha 4 high

Corrected grading following fixing trac's priority list..

comment:13 Changed 10 years ago by mark1

  • Owner mark1 deleted

comment:14 Changed 10 years ago by emca

  • Owner set to emca

Starting LIDAR processing

comment:15 Changed 10 years ago by emca

LIDAR processing is complete and project is now ready for delivery checking.
Delivery directory created at
~airborne/workspace/GB08_08-2009_075b_Wytham_Woods/delivery

comment:16 Changed 10 years ago by mark1

  • Owner changed from emca to mark1
  • Status changed from new to assigned

comment:17 Changed 10 years ago by mark1

  • Owner changed from mark1 to emca

Delivery checked and looks good. Data viewed in GIS and lines up OK. Ready to deliver.

Data has not been rsynced over.

comment:18 Changed 10 years ago by emca

LIDAR data sent to PI 01/12/2009

comment:19 Changed 10 years ago by emca

LIDAR data has been rsynced to
/users/rsg/arsf/arsf_data/2009/flight_data/uk/GB08_08-2009_075b_Wytham_Woods

comment:20 Changed 9 years ago by mark1

Photography delivered on hard disk 05/01/2010

comment:21 Changed 9 years ago by mark1

Photography delivery rsynced to

/users/rsg/arsf/arsf_data/2009/flight_data/uk/GB08_08-2009_075b_Wytham_Woods

comment:22 Changed 9 years ago by emca

  • Component changed from Processing: general to Archiving
  • Owner changed from emca to benj

comment:23 Changed 9 years ago by knpa

  • Owner changed from benj to knpa

Archived by DTG (12/01/10).
Preparing for NEODC.

comment:24 Changed 9 years ago by knpa

Tarball created and NEODC informed ready for transfer.

comment:25 Changed 9 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed

Transferred to NEODC.
Closing ticket.

comment:26 Changed 6 years ago by stgo

SCT details from config file.

E H
1-0.055-0.09
2-0.055-0.12
3-0.055-0.055
4-0.055-0.04

comment:27 Changed 2 years ago by asm

Archiving

Reprocessed hyperspectral has been archived and it is available from CEDA at http://browse.ceda.ac.uk/browse/neodc/arsf/2009/GB08_08/GB08_08-2009_075b_Wytham_Woods/hyperspectral-reprocessed

Plymouth 11th May 2017

Last edited 2 years ago by asm (previous) (diff)
Note: See TracTickets for help on using tickets.